The radius R of collapsing spinning star drops to 1/3 its initial value wheres the mass M remains unchanged. The star spins as a rigid body and its rotational inertia I =2/5 MR^2. The ratio of the new rotational kinetic energy to the initial rotational kinetic energy is
1
Expert's answer
2019-06-12T15:24:26-0400

From the conservation of angular momentum:


"I\omega=I'\omega'"

"\frac{2}{5}MR^2\omega=\frac{1}{3^2}\frac{2}{5}MR^2\omega'"

"\omega'=9\omega"

"I'=\frac{1}{9}I"

The ratio of the new rotational kinetic energy to the initial rotational kinetic energy is


"\frac{0.5I'\omega'^2}{0.5I\omega^2}=9^2\frac{1}{9}=9"